A car travels at a constant speed around a circular track whose radius is 3.37 km. The car goes once around the track in 226 s. What is the magnitude of the centripetal acceleration of the car?

I get 1.88 m/s^2
Answer #1

here,

the radius of track , r = 3.37 km = 3370 m

time period , T = 226 s

the angular speed , w = 2 *pi /T

w = 2 *pi /226 rad/s

w = 0.02779 rad/s

the magnitude of centripetal acceleration , ac = w^2 * r

ac = 0.02779^2 * 3370 m/s^2

ac = 2.6 m/s^2